OpenStudy (anonymous):

Find the sixth term of a geometric sequence with t5 = 24 and t8 = 3.

OpenStudy (welshfella):

each term is derived from the previous by multiplying by the common ratio r so t8 / t5 = r^3 so what id r?

OpenStudy (welshfella):

24/3 = r^3

OpenStudy (anonymous):

8

OpenStudy (welshfella):

r^3 = 8 not r

OpenStudy (anonymous):

so do i devide by 3

OpenStudy (welshfella):

no you find the cube root of 8 r^3 = r * r * r

OpenStudy (welshfella):

what number when multiplied by itself 3 times = 8 ?

OpenStudy (anonymous):

um like 2.6 maybe

OpenStudy (welshfella):

2.6 * 2.6 * 2.6 = 17.58

OpenStudy (welshfella):

- i think you are referring to the square root we want the cube root

OpenStudy (anonymous):

2.7?

OpenStudy (welshfella):

no its 2 2*2*2 = 8

OpenStudy (anonymous):

oh im dumb omg

OpenStudy (anonymous):

sorry

OpenStudy (welshfella):

lol - ok dont fret!!

OpenStudy (welshfella):

- don't think you are the only one to make a kistake I've just realised I've made one right from the beginning!

OpenStudy (anonymous):

i was thinking 2.7x3....its okay lol

OpenStudy (welshfella):

t8 = 3 and t5 = 24 so r^3 = 3/24 not 24/3 so r = 1/8 not 8

OpenStudy (anonymous):

ohhh so what do i do from here

OpenStudy (welshfella):

ok so r = 1/2 now each term is derived from the last by multiplying by 1/2 so 5th term = 24 so 6th term = 24 * 1/2 = 12

OpenStudy (welshfella):

* r^3 = 1/8 so r = 1/2

OpenStudy (anonymous):

so how did we get the 1/2

OpenStudy (welshfella):

6th term = 12

OpenStudy (welshfella):

cube root of 1/8 is 1/2

OpenStudy (anonymous):

okay and is their a way to find the cube rute on a calculater

OpenStudy (welshfella):

1/2 * 1/2 * 1/2 = 1/8

OpenStudy (welshfella):

yes with mine you have a power key which is ^ now cube root is the same as power 1/3 so I would enter 1/8 ^ 1/3

OpenStudy (welshfella):

though some calculators have a cube root key

OpenStudy (anonymous):

okay thanks..Find the three geometric means between 2 and 162. ........do u know how to do these kinds

OpenStudy (anonymous):

if not it is okay u have helped me so much

OpenStudy (welshfella):

lol I've forgotten how to do those ...

OpenStudy (anonymous):

okay thanks anyways

OpenStudy (welshfella):

if no one can help try google

OpenStudy (welshfella):

yw